Circular No. 8796

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 A3 (PETRIEW) Below are B. G. Marsden’s epoch-2001 orbital elements from the 2001-

2007 observations (cf. IAUC 8795):

Epoch = 2001 Sept. 8.0 TT

T = 2001 Sept. 1.9221 TT  ω = 181.^{o}9012
e = 0.696237  Ω = 214.1070 }2000.0
q = 0.945691 AU  i = 13.9444

a = 3.113256 AU n^{o} = 0.1794246 P = 5.49 years

COMET P/2007 A2 (CHRISTENSEN) = 2006 WY_{182} This comet (cf. IAUC 8794) has been linked with 2006 WY_{182} (cf. MPS

190906); elliptical orbital elements from MPEC 2007-A47:

T = 2007 Jan. 17.5046 TT  ω = 186.^{o}8829
e = 0.558240  Ω = 301.0412 }2000.0
q = 2.795797 AU  i = 26.4771
a = 6.328771 AU n^{o} = 0.0619049 P = 15.92 years

COMET P/2006 Y2 (GIBBS) This comet (cf. IAUC 8787) has also been found to be of short orbital

period, the following elements taken from MPEC 2007-A52:

T = 2006 Dec. 28.7559 TT  ω = 26.^{o}3029
e = 0.587452  Ω = 113.3627 }2000.0
q = 1.251352 AU  i = 14.4737
a = 3.033225 AU n^{o} = 0.1865723 P = 5.28 years

COMET C/2006 P1 (McNAUGHT) Additional visual total-magnitude estimates made in daylight, via com-

parison with Venus, many observers noting a tail ≤ 1^{o} long: Jan. 13.10 UT, −5: (S. J. O’Meara, Mauna Kea, HI, U.S.A., naked eye); 13.56, −6: (A. Amorim, Florianopolis, Brazil, naked eye and sunglasses); 13.75, −5.2: (R. A. Keen, Mt. Thorodin, CO, U.S.A., naked eye); 13.91, −5.0: (C. S. Morris, Fillmore, CA, U.S.A., 20×80 binoculars); 13.92, −6: (O’Meara, Volcano, HI); 14.47, −5.0: (D. J. Frew, Perth Observatory, 8×42 binoculars); 14.55, −5.5: (K. Hornoch, Lelekovice, Czech Republic, naked eye).

2007 January 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green